# Facial Recognition Leveraging Generative Adversarial Networks

### v1

@@ 1. Introduction
Face recognition systems have reached remarkable accuracy levels when trained on large-scale datasets, yet their performance degrades significantly in data-scarce scenarios a common challenge in specialized applications such as medical diagnostics or forensic analysis (wang2021generative, ). While Generative Adversarial Networks (GANs) have shown promise for small-sample augmentation, current approaches suffer from two critical limitations: (1) generated images often lack discriminative facial features crucial for recognition (2) existing frameworks are not optimized for integration with modern face recognition architectures.
@@ 2. Background
The current research in face recognition technology mainly includes the development of deep learning methods the establishment of large-scale datasets, multimodal fusion, cross-domain face recognition(niu2024unveiling, ), robustness and privacy protection, and other aspects. With the rapid development of deep learning technology, face recognition methods based on deep neural networks have achieved great success, especially the application of convolutional neural networks in face recognition tasks (9724923, ).
For a face recognition model to achieve better recognition results, it is usually necessary to focus on the following two aspects:  (1) data quality and quantity  (2) model architecture and parameter design. ResNet (Residual Network), FaceNet, DenseNet (Dense Connected Network), and SENet (Attention Mechanism Network) (mehta2022three, ), are some of the widely recognized as better face recognition model architectures. The above have achieved excellent performance on large-scale datasets celebA and VGGFcae, among others. However, data acquisition and labeling require a lot of resources and time for face recognition tasks(kong2024characterizing, ).
In some specific complex environments, the inability to obtain enough sample data to train the model often leads to model overfitting, which reduces its generalizability. And small sample data augmentation is a commonly used approach to address the lack of data (liu2023deep, ). GAN makes it possible to provide high-quality sample data for face recognition models by virtue of its dynamic gaming as well as its powerful image generation capability . Meanwhile, by utilizing the dynamic game characteristics of GAN and using the face recognition model as the discriminator of GAN, the degree of difference between the fake face image generated by the generator and the real face image can be further improved, which in turn improves the generating ability of the generator and the quality of the generated image  (alqahtani2021applications, ). In this way, the GAN can generate more realistic and high-quality face images, thus improving the accuracy and robustness of the face recognition system.
@@ 2.1. GAN model
GAN is a deep learning model that consists of two main components: Generator and Discriminator. The core of GAN is that the two components compete with each other to improve performance through constant gaming, so as to achieve the goal of generating high-quality data. The Generator is responsible for generating synthetic data similar to the real data. It receives a random noise vector as input and maps it to the output space through a series of transformations to generate fake data (li2021clue, ). The discriminator is responsible for distinguishing between real data and synthetic data generated by the generator, it receives data (real or synthetic) as input and outputs a probability indicating the probability that the input data is real. During the training process, the generator and discriminator are trained alternately so that the synthetic data generated by the generator can be more and more realistic, while the discriminator can more accurately distinguish between real data and synthetic data (alqahtani2021applications, ).
The loss function of the GAN model consists of two parts: the generator’s loss and the discriminator’s loss. The generator’s loss function is designed to make the generated synthetic data closer to the real data, while the discriminator’s loss function is designed to be more accurate in distinguishing between real and synthetic data (li2024stateguard, ). The GAN model has been widely used in several fields, including image generation, image editing, video generation, speech synthesis, and so on. It not only generates realistic images but also can be used for tasks such as data enhancement and domain adaptation (mao2024scla, ).
Although GAN models have achieved many successful applications, they still face some challenges, such as pattern collapse and unstable training. To solve these problems, researchers have proposed many improvements such as Wasserstein GAN, Self-Attention GAN (hwang2023adversarial, ).
@@ 2.2. Face Recognition Technology
Face recognition technology is a biometric identification technology based on face feature information. This chapter summarizes traditional face recognition methods and their advantages and disadvantages and focuses on deep learning-based face recognition technology (bu2025smartbugbert, ). Among them, GAN-based face recognition research has achieved remarkable results in recent years, which provides an important theoretical foundation for the work of this paper.
Face recognition techniques can be divided into two categories: face recognition based on traditional methods and face recognition based on deep learning (li2020review, ). Traditional methods mainly rely on manually designed feature extractors and machine learning algorithms, while deep learning methods utilize deep neural networks to automatically learn feature representations and classification rules (wang2024smart, ).
@@ 2.2.1. Traditional face recognition methods and their advantages and disadvantages
The eigenfaces method is a traditional face recognition technique based on principal component analysis (PCA). The face image is represented as vectors in a high-dimensional feature space, and these vectors are downscaled and feature extracted using PCA to obtain a set of principal components or feature vectors called “eigenfaces”. By training the classifier, the featured face is used for face recognition (adjabi2020past, ).
@@
In face recognition, LDA achieves feature extraction by maximizing the interclass distance and minimizing the intraclass distance.LDA projects the data into a low-dimensional space so that the distance between samples of the same class is as small as possible and the distance between samples of different classes is as large as possible (alqahtani2021applications, ).
@@
The advantages of these traditional methods are that they are simple to implement, easy to understand, and suitable for small-scale datasets and simple scenes (li2025scalm, ). However, traditional face recognition methods such as eigenface methods, linear discriminant analysis, and local binary patterns have some limitations and drawbacks in practical applications, such as sensitivity to factors such as illumination, pose, expression, etc., as well as limited ability to deal with problems such as imbalance in the sample categories, image noise and scale variations (yan2021enhanced, ).
@@ 2.2.2. Deep learning based face recognition techniques
CNN is able to extract high-level abstract features from the original image through multi-layer convolution and pooling operations, which include edges, textures, shapes, etc., and help to recognize different faces. The deep structure of CNN ( ResNet, VGG) has been applied to the face recognition task, and these models have been trained on large-scale datasets and are able to learn richer and more abstract feature representations (wu2021mtcnn, ).
@@
ResNet refers to the method of face feature learning and recognition using residual networks.ResNet is a deep neural network structure proposed by Microsoft Research, which solves the problem of gradient vanishing and gradient explosion that occurs during the training of deep neural networks by introducing residual blocks, making it possible to train very deep network models (mehta2022three, ).
@@
A deep learning-based face recognition model that uses a training method called Triplet Loss to achieve efficient face verification and recognition by mapping the face images of the same person into a similar embedding space while mapping the face images of different people into a more distant embedding space. The structure of FaceNet (schroff2015facenet, ) is shown in Fig 2.
@@
Deep learning-based face recognition techniques, although in have high accuracy, can be trained end-to-end, learn feature representation and classification directly from raw data, reduce the process of manually processing data as well and be able to automatically discover and utilize local and global information in images (li2024detecting, ). However, it usually requires a large amount of labeled data for training, especially in the field of face recognition, which requires large-scale face image datasets. Recent studies have shown that blockchain-based decentralized data markets can help alleviate this data scarcity issue while preserving privacy (10707457, ). Deep learning models are susceptible to adversarial attacks (kammoun2022generative, ), which can lead to incorrect model outputs through small perturbations, posing challenges to model security. The immutable audit trail provided by smart contracts (bu2025enhancingsmartcontractvulnerability, ) offers potential solutions for detecting and preventing such attacks in facial recognition systems.
@@ 2.2.3. Research on GAN-based face recognition
The current research on GAN-based face recognition mainly involves the following four aspects: (1) using GAN to generate high-fidelity and high-resolution face images to solve the problem of data scarcity and sample imbalance, and to improve the richness and diversity of training data, though recent work by (li2021hybrid, ) suggests blockchain-based verification could further ensure generated data authenticity. (2) generate adversarial samples using GAN to expand the training dataset and improve the generalization ability and robustness of face recognition models, with security considerations aligning with the systemic risk framework in (liu2025sok, ). (3) using GAN for cross-domain face recognition to achieve feature migration and knowledge migration between different datasets to improve the generalization performance and adaptability of the model. (4) using GAN to generate adversarial samples, study the mechanism and defense methods of adversarial attacks, and improve the security and anti-interference ability of face recognition systems(cui2019image, ), with (li2017discovering, ) demonstrating similar vulnerability patterns in other machine learning domains and providing cross-domain security analysis methodologies.
The advantages of the application of the GAN model in the field of images are very obvious, but it also has its own disadvantages. It is pointed out in the literature that the GAN model is prone to gradient disappearance, explosion, and training instability during training, issues that parallel the smart contract vulnerabilities identified in (hybrid_analysis) and systematically categorized in (liu2025sok, )’s taxonomy of decentralized system failures. At the same time, the literature(kammoun2022generative, ) also mentioned the problems related to the GAN model, and proposed the use of residual network and residual block to solve the problem of gradient disappearance gradient explosion and training crash encountered in the training process of GAN model.
@@ 3.1. Overall framework of GAN-based face recognition
In GAN-based face recognition, the residual block technique can be applied to improve the structure of the generator and the discriminator to improve the performance and stability of the model. By introducing residual connectivity, the gradient can be better propagated and the convergence of the network accelerated, while the problem of gradient vanishing during training is reduced. This makes the GAN-based face recognition model easier to optimize and train while improving the quality and realism of the generated face images (banerjee2018lr, ).
@@ 3.2. Generator
When the initial face image is input, the input image undergoes a convolution operation through a 9×9 convolutional layer, this convolutional layer extracts the basic features such as the edges and texture of the input face, and then the feature map undergoes two 3×3 convolutional layers, each of which reduces the size of the feature map while increasing the depth of the features. After passing through two convolutional layers, the feature map is normalized by three spectral normalization layers. Finally, the feature map passes through six residual blocks. Higher-level information features such as contours, details, and structures of the face are learned and the encoder structure is shown in Fig 4.
@@ 3.4. Discriminator
The discriminator uses the FaceNet model based on the Inception Resnet V1 architecture, which combines the advantages of the Inception module and the Residual Connection. The Inception module processes the input data in parallel with multiple convolutional kernels of different sizes to capture features at different scales, employing a multi-scale analysis approach similar to the opcode vectorization strategy used in smart contract security analysis (zou2025malicious, ). The Residual Connection, on the other hand, helps to solve the problem of gradient vanishing and representation bottleneck in deep networks, allowing the network to stack deeper without performance degradation (9, ). When this structure is applied to face recognition, FaceNet is able to extract richer and more robust face features. These features are more robust to changes in illumination, expression, and posture, thus improving the accuracy of face recognition.
@@
The FaceNet model works by converting face images into feature vectors and mapping these feature vectors into a multidimensional space. In this multidimensional space, the feature vectors corresponding to the face images of the same person are close together, while the feature vectors corresponding to the face images of different people are farther apart. By calculating the distance between these feature vectors in Euclidean space, the similarity between faces can be evaluated, with stability verification principles adapted from the pattern consistency detection in (zou2025malicious, ). If the distance between the feature vectors of two face images is less than a certain threshold, they are considered to be from the same person; otherwise, they are considered to be from different people. This method can effectively distinguish faces and achieve face recognition. The Euclidean spatial distance formula is shown below, where d represents the distance, and $x_{1m}$,$x_{2m}$ represent the feature vectors of the two images.
@@ 4.4.2. Experimental method and procedure
Fig $1$ demonstrates the process of synthesizing adversarial samples by the synthesizer. Where Fig 11 shows the adversarial perturbation, Fig 11 shows the salient feature extraction, and Fig 11 shows the generated adversarial sample. The salient map reflects the importance of each part of the face image, the higher the importance of the region in the face image corresponds to a larger value of the salient map, and the lower the importance of the region corresponds to a smaller value of the salient map. From the figure, it can be seen that the saliency map generated by this method can well determine different perturbation regions according to different inputs, and generate different weights for the antiperturbation according to the relative importance of each part of the input image, thus improving the image quality of the antiperturbation samples. With the input of about 5000 face images from the initial AR face library dataset, after the random feature perturbations are added by the GAN model generator and the salient region extraction module in this paper, finally about 500000 (0.5M) sample images are generated for the training of the discriminator.
@@ 4.5. Model Implementation
The train-GAN file implements the training process with security monitoring adapted from (li2024cobra, )’s vulnerability detection framework, including: (1) Loading the data with integrity checks (2) Model parameter validation (3) Training process recording with anomaly detection
@@ 6. Conclusion
Through several months of related literature reading and experiments, we realized the construction of the GAN model using the residual network and solved the problems of gradient disappearance and explosion encountered in the training process of the GAN model. At the same time contend with the face recognition model under complex conditions, the training samples are difficult to obtain, and the lack of sufficient samples leads to poor model robustness and generalization ability. In this paper, the possibility of small sample data enhancement using the GAN model. The significant feature extraction module is added to the GAN model realized by the residual network to extract the significant features of the face and randomly add perturbations to it. Small sample data enhancement is successfully performed to expand the face recognition model training set. Using the dynamic game characteristics of the GAN model, the FaceNet model based on the Inception Resnet V1 structure is used as the discriminator of the GAN, which further improves the recognition ability of the face recognition model. Finally, the feasibility of the work done in this paper is verified through validation experiments and comparison experiments.

### v2

@@ 1. Introduction
Face recognition systems have reached remarkable accuracy levels when trained on large-scale datasets, yet their performance degrades significantly in data-scarce scenarios a common challenge in specialized applications such as medical diagnostics or forensic analysis (Wang et al., 2021). While Generative Adversarial Networks (GANs) have shown promise for small-sample augmentation, current approaches suffer from two critical limitations: (1) generated images often lack discriminative facial features crucial for recognition (2) existing frameworks are not optimized for integration with modern face recognition architectures.
@@ 2. Background
The current research in face recognition technology mainly includes the development of deep learning methods the establishment of large-scale datasets, multimodal fusion, cross-domain face recognition(Niu et al., 2024), robustness and privacy protection, and other aspects. With the rapid development of deep learning technology, face recognition methods based on deep neural networks have achieved great success, especially the application of convolutional neural networks in face recognition tasks (Wang and Guo, 2021).
For a face recognition model to achieve better recognition results, it is usually necessary to focus on the following two aspects: (1) data quality and quantity (2) model architecture and parameter design. ResNet (Residual Network), FaceNet, DenseNet (Dense Connected Network), and SENet (Attention Mechanism Network) (Mehta et al., 2022; Jiao et al., 2024; Kumar et al., 2024; Wei et al., 2024), are some of the widely recognized as better face recognition model architectures. The above have achieved excellent performance on large-scale datasets celebA and VGGFcae, among others. However, data acquisition and labeling require a lot of resources and time for face recognition tasks(Kong et al., 2024; Zhu et al., 2024).
In some specific complex environments, the inability to obtain enough sample data to train the model often leads to model overfitting, which reduces its generalizability. And small sample data augmentation is a commonly used approach to address the lack of data (Liu et al., 2023). GAN makes it possible to provide high-quality sample data for face recognition models by virtue of its dynamic gaming as well as its powerful image generation capability . Meanwhile, by utilizing the dynamic game characteristics of GAN and using the face recognition model as the discriminator of GAN, the degree of difference between the fake face image generated by the generator and the real face image can be further improved, which in turn improves the generating ability of the generator and the quality of the generated image (Alqahtani et al., 2021). In this way, the GAN can generate more realistic and high-quality face images, thus improving the accuracy and robustness of the face recognition system.
@@ 2.1. GAN model
GAN is a deep learning model that consists of two main components: Generator and Discriminator. The core of GAN is that the two components compete with each other to improve performance through constant gaming, so as to achieve the goal of generating high-quality data. The Generator is responsible for generating synthetic data similar to the real data. It receives a random noise vector as input and maps it to the output space through a series of transformations to generate fake data (Li et al., 2021). The discriminator is responsible for distinguishing between real data and synthetic data generated by the generator, it receives data (real or synthetic) as input and outputs a probability indicating the probability that the input data is real. During the training process, the generator and discriminator are trained alternately so that the synthetic data generated by the generator can be more and more realistic, while the discriminator can more accurately distinguish between real data and synthetic data (Alqahtani et al., 2021).
The loss function of the GAN model consists of two parts: the generator’s loss and the discriminator’s loss. The generator’s loss function is designed to make the generated synthetic data closer to the real data, while the discriminator’s loss function is designed to be more accurate in distinguishing between real and synthetic data (Li et al., 2024d). The GAN model has been widely used in several fields, including image generation, image editing, video generation, speech synthesis, and so on. It not only generates realistic images but also can be used for tasks such as data enhancement and domain adaptation (Mao et al., 2024).
Although GAN models have achieved many successful applications, they still face some challenges, such as pattern collapse and unstable training. To solve these problems, researchers have proposed many improvements such as Wasserstein GAN, Self-Attention GAN (Hwang et al., 2023).
@@ 2.2. Face Recognition Technology
Face recognition technology is a biometric identification technology based on face feature information. This chapter summarizes traditional face recognition methods and their advantages and disadvantages and focuses on deep learning-based face recognition technology (Bu et al., 2025b). Among them, GAN-based face recognition research has achieved remarkable results in recent years, which provides an important theoretical foundation for the work of this paper.
Face recognition techniques can be divided into two categories: face recognition based on traditional methods and face recognition based on deep learning (Li et al., 2020). Traditional methods mainly rely on manually designed feature extractors and machine learning algorithms, while deep learning methods utilize deep neural networks to automatically learn feature representations and classification rules (Wang et al., 2024b).
@@ 2.2.1. Traditional face recognition methods and their advantages and disadvantages
The eigenfaces method is a traditional face recognition technique based on principal component analysis (PCA). The face image is represented as vectors in a high-dimensional feature space, and these vectors are downscaled and feature extracted using PCA to obtain a set of principal components or feature vectors called “eigenfaces”. By training the classifier, the featured face is used for face recognition (Adjabi et al., 2020).
@@
In face recognition, LDA achieves feature extraction by maximizing the interclass distance and minimizing the intraclass distance.LDA projects the data into a low-dimensional space so that the distance between samples of the same class is as small as possible and the distance between samples of different classes is as large as possible (Alqahtani et al., 2021).
@@
The advantages of these traditional methods are that they are simple to implement, easy to understand, and suitable for small-scale datasets and simple scenes (Li et al., 2025b). However, traditional face recognition methods such as eigenface methods, linear discriminant analysis, and local binary patterns have some limitations and drawbacks in practical applications, such as sensitivity to factors such as illumination, pose, expression, etc., as well as limited ability to deal with problems such as imbalance in the sample categories, image noise and scale variations (Yan et al., 2021).
@@ 2.2.2. Deep learning based face recognition techniques
CNN is able to extract high-level abstract features from the original image through multi-layer convolution and pooling operations, which include edges, textures, shapes, etc., and help to recognize different faces. The deep structure of CNN ( ResNet, VGG) has been applied to the face recognition task, and these models have been trained on large-scale datasets and are able to learn richer and more abstract feature representations (Wu and Zhang, 2021).
@@
ResNet refers to the method of face feature learning and recognition using residual networks.ResNet is a deep neural network structure proposed by Microsoft Research, which solves the problem of gradient vanishing and gradient explosion that occurs during the training of deep neural networks by introducing residual blocks, making it possible to train very deep network models (Mehta et al., 2022).
@@
A deep learning-based face recognition model that uses a training method called Triplet Loss to achieve efficient face verification and recognition by mapping the face images of the same person into a similar embedding space while mapping the face images of different people into a more distant embedding space. The structure of FaceNet (Schroff et al., 2015) is shown in Fig 2.
@@
Deep learning-based face recognition techniques, although in have high accuracy, can be trained end-to-end, learn feature representation and classification directly from raw data, reduce the process of manually processing data as well and be able to automatically discover and utilize local and global information in images (Li et al., 2024b). However, it usually requires a large amount of labeled data for training, especially in the field of face recognition, which requires large-scale face image datasets. Recent studies have shown that blockchain-based decentralized data markets can help alleviate this data scarcity issue while preserving privacy (Liu et al., 2024). Deep learning models are susceptible to adversarial attacks (Kammoun et al., 2022), which can lead to incorrect model outputs through small perturbations, posing challenges to model security. The immutable audit trail provided by smart contracts (Bu et al., 2025a) offers potential solutions for detecting and preventing such attacks in facial recognition systems.
@@ 2.2.3. Research on GAN-based face recognition
The current research on GAN-based face recognition mainly involves the following four aspects: (1) using GAN to generate high-fidelity and high-resolution face images to solve the problem of data scarcity and sample imbalance, and to improve the richness and diversity of training data, though recent work by (Li and others, 2021) suggests blockchain-based verification could further ensure generated data authenticity. (2) generate adversarial samples using GAN to expand the training dataset and improve the generalization ability and robustness of face recognition models, with security considerations aligning with the systemic risk framework in (Liu and Li, 2025). (3) using GAN for cross-domain face recognition to achieve feature migration and knowledge migration between different datasets to improve the generalization performance and adaptability of the model. (4) using GAN to generate adversarial samples, study the mechanism and defense methods of adversarial attacks, and improve the security and anti-interference ability of face recognition systems(Cui et al., 2019), with (Li et al., 2017) demonstrating similar vulnerability patterns in other machine learning domains and providing cross-domain security analysis methodologies.
The advantages of the application of the GAN model in the field of images are very obvious, but it also has its own disadvantages. It is pointed out in the literature that the GAN model is prone to gradient disappearance, explosion, and training instability during training, issues that parallel the smart contract vulnerabilities identified in (hybrid_analysis; Wang and Zhao, 2024; Li et al., 2025a, 2024c) and systematically categorized in (Liu and Li, 2025)’s taxonomy of decentralized system failures. At the same time, the literature(Kammoun et al., 2022; Wang et al., 2024a, c) also mentioned the problems related to the GAN model, and proposed the use of residual network and residual block to solve the problem of gradient disappearance gradient explosion and training crash encountered in the training process of GAN model.
@@ 3.1. Overall framework of GAN-based face recognition
In GAN-based face recognition, the residual block technique can be applied to improve the structure of the generator and the discriminator to improve the performance and stability of the model. By introducing residual connectivity, the gradient can be better propagated and the convergence of the network accelerated, while the problem of gradient vanishing during training is reduced (Wu et al., 2025; Xiao et al., 2025; Cai et al., 2024; Han et al., 2026). This makes the GAN-based face recognition model easier to optimize and train while improving the quality and realism of the generated face images (Banerjee and Das, 2018).
@@ 3.2. Generator
When the initial face image is input, the input image undergoes a convolution operation through a 9×9 convolutional layer, this convolutional layer extracts the basic features such as the edges and texture of the input face, and then the feature map undergoes two 3×3 convolutional layers, each of which reduces the size of the feature map while increasing the depth of the features. After passing through two convolutional layers, the feature map is normalized by three spectral normalization layers. Finally, the feature map passes through six residual blocks. Higher-level information features such as contours, details, and structures of the face are learned and the encoder structure is shown in Fig 4. The preservation of these fundamental structures is essential for maintaining semantic consistency during generation (Prifti et al., 2024; Su et al., 2025; Wei et al., 2025; Zhuang et al., 2021).
@@ 3.4. Discriminator
The discriminator uses the FaceNet model based on the Inception Resnet V1 architecture, which combines the advantages of the Inception module and the Residual Connection. The Inception module processes the input data in parallel with multiple convolutional kernels of different sizes to capture features at different scales, employing a multi-scale analysis approach similar to the opcode vectorization strategy used in smart contract security analysis (Zou et al., 2025). The Residual Connection, on the other hand, helps to solve the problem of gradient vanishing and representation bottleneck in deep networks, allowing the network to stack deeper without performance degradation (Tong et al., 2024). When this structure is applied to face recognition, FaceNet is able to extract richer and more robust face features. These features are more robust to changes in illumination, expression, and posture, thus improving the accuracy of face recognition.
@@
The FaceNet model works by converting face images into feature vectors and mapping these feature vectors into a multidimensional space. In this multidimensional space, the feature vectors corresponding to the face images of the same person are close together, while the feature vectors corresponding to the face images of different people are farther apart. By calculating the distance between these feature vectors in Euclidean space, the similarity between faces can be evaluated, with stability verification principles adapted from the pattern consistency detection in (Zou et al., 2025). If the distance between the feature vectors of two face images is less than a certain threshold, they are considered to be from the same person; otherwise, they are considered to be from different people. This method can effectively distinguish faces and achieve face recognition. The Euclidean spatial distance formula is shown below, where d represents the distance, and $x_{1m}$$x_{2m}$ represent the feature vectors of the two images.
@@ 4.4.2. Experimental method and procedure
Fig $11$ demonstrates the process of synthesizing adversarial samples by the synthesizer. Where Fig 11 shows the adversarial perturbation, Fig 11 shows the salient feature extraction, and Fig 11 shows the generated adversarial sample. The salient map reflects the importance of each part of the face image, the higher the importance of the region in the face image corresponds to a larger value of the salient map, and the lower the importance of the region corresponds to a smaller value of the salient map. From the figure, it can be seen that the saliency map generated by this method can well determine different perturbation regions according to different inputs, and generate different weights for the antiperturbation according to the relative importance of each part of the input image, thus improving the image quality of the antiperturbation samples. With the input of about 5000 face images from the initial AR face library dataset, after the random feature perturbations are added by the GAN model generator and the salient region extraction module in this paper, finally about 500000 (0.5M) sample images are generated for the training of the discriminator.
@@ 4.5. Model Implementation
The train-GAN file implements the training process with security monitoring adapted from (Li et al., 2024a)’s vulnerability detection framework, including: (1) Loading the data with integrity checks (2) Model parameter validation (3) Training process recording with anomaly detection
@@ 6. Conclusion
Through several months of related literature reading and experiments, we realized the construction of the GAN model using the residual network and solved the problems of gradient disappearance and explosion encountered in the training process of the GAN model. At the same time contend with the face recognition model under complex conditions, the training samples are difficult to obtain, and the lack of sufficient samples leads to poor model robustness and generalization ability. In this paper, the possibility of small sample data enhancement using the GAN model. The significant feature extraction module is added to the GAN model realized by the residual network to extract the significant features of the face and randomly add perturbations to it. Small sample data enhancement is successfully performed to expand the face recognition model training set. Using the dynamic game characteristics of the GAN model, the FaceNet model based on the Inception Resnet V1 structure is used as the discriminator of the GAN, which further improves the recognition ability of the face recognition model. Finally, the feasibility of the work done in this paper is verified through validation experiments and comparison experiments (Aguilar et al., 2024; Arceri et al., 2024; Ayub et al., 2024; Chen et al., 2024; Grossman et al., 2024).